OpenStudy (anonymous):

Simplify the expression. 3 + 22 · 4 – 1 I got 18, is that right?!

OpenStudy (anonymous):

I think it should be 90 22*4 = 88 +3 = 91 -1 = 90

OpenStudy (anonymous):

oh yeah the 22 is just a 2 with an Exponent of 2

OpenStudy (anonymous):

Then it would be 4*4 = 16 + 3 = 19 - 1= 18

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

:)